How much force must be exerted through the wire rope that moves the car in an elevator given that the car masses 501kg and there are 7 people average mass 75 kg and that the friction in the guide along the shaft average 205N. when
a) the car is accelerating upward at 1.00m/s^2?
b) the car is moving upward and slowing down at the equivalent rate?
1
Expert's answer
2019-11-24T17:07:26-0500

Point y-axis up, then the force of friction and force of gravity both look down (negative).

a) Here the acceleration is positive (it looks up). According to Newton's second law:


"Ma=-f-Mg+F,\\
F=Ma+f+Mg=M(g+a)+f=\\=(501+7\cdot75)(9.8+1)+205=11285.8\text{ N}."


b) In this case everything's the same, but the acceleration "a" is directed downward:


"-Ma=-f-Mg+F,\\
F=f+Mg-Ma=M(g+a)+f=\\=205+(501+7\cdot75)(9.8-1)=92338\text{ N}."